OpenStudy (anonymous):

If the ratio of the heights of two similar cylinders is 1/4, what is the ratio of their volumes?

OpenStudy (mayankdevnani):

volume of cylinder=pi*r^2*h

OpenStudy (mayankdevnani):

ok @janett77

OpenStudy (mayankdevnani):

got it?

OpenStudy (mayankdevnani):

\[\frac{h1}{h2}=\frac{1}{4}\]

OpenStudy (mayankdevnani):

\[\frac{\Pi \times r^2\times1}{\Pi \times r^2\times4}\]

OpenStudy (anonymous):

is the answer 1/16

OpenStudy (mayankdevnani):

no answer is 1/4

OpenStudy (mayankdevnani):

ok @janett77

OpenStudy (anonymous):

why is it 1/4

OpenStudy (mayankdevnani):

because